OpenStudy (anonymous):

What is the value of n in the equation 12n + 8 = 11? Please submit your answer as a reduced fraction.

OpenStudy (mathmale):

In other words, solve this equation for n. Subtract 8 from both sides, to isolate 12n on the left side. Simplify the right side. Write the equation as 12n = (right side). Solve for n by dividing this equation by 12.

OpenStudy (anonymous):

wait so like this

OpenStudy (anonymous):

12n+8=11 -8 -8

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

You're on the right track, but need to finish the work you've started. Hint: 8-8=0. Finish this, please: 12n+8=11 -8 -8 ------ ---

OpenStudy (anonymous):

8-8=0 11-8=3

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

ok so its 12x-9+8=11

OpenStudy (anonymous):

sorry 12+-9+8=11

OpenStudy (mathmale):

12n+8=11 -8 -8 ------ --- 12n = 3 (not 12x-9+8=11) if 12n = 3, solve for n.

OpenStudy (anonymous):

oh ok so whtas the next step @mathmale

OpenStudy (mathmale):

Please don't be afraid to take risks. I'd like for YOU to make the next move. As before, 12n = 3. Solve this for n. How have you done this in the past?

OpenStudy (anonymous):

n=4

OpenStudy (mathmale):

Take 12n = 3 and divide both sides by 12. Note: if n=4, then 12(4) would not equal 3. So n=4 could not be the correct answer. Divide both sides of 12n=3 by 12.

OpenStudy (anonymous):

3/12=0.25

OpenStudy (mathmale):

That's better. There's one other way in which to write your answer: 1/4. So: n=1/4.

OpenStudy (mathmale):

Please review what we've done here, so that you'll know better what to do next time.

OpenStudy (anonymous):

thanks! @mathmale

OpenStudy (mathmale):

You're welcome!! Thanks for the medal.